# Introspective Planning: Guiding Language-Enabled Agents to Refine Their Own Uncertainty

### 09 Feb 2024 v1

@@ 2 Introspective Planning
The fundamental aim of introspective planning is to guide LLMs automatically to reason about task uncertainty and assess the feasibility of candidate actions. While this guidance can take various forms, our implementation here is based on distilling example reasoning processes from a knowledge base to inform the LLM via in-context learning. Problem Formulation. Similar to (Ren et al., 2023), we cast LLM-based planning as multiple-choice question answering (MCQA). Given a task statement $d_{i}$ and the observation $o_{i}$, the LLM planner first generates a set of candidate plans $\mathcal{C}_{i}$, each assigned a unique letter label from $\mathcal{Y}\coloneqq\{\textrm{`}A\textrm{'},\textrm{`}B\textrm{'},\textrm{`}C\textrm{'},\ldots\}$ ’ , … }. The planner then predicts $\hat{y}_{i}\in\mathcal{Y}$, aiming to match the unknown true user intent $z_{i}$. For example, consider the stated task $d_{i}$ring me that soda” with the observation $o_{i}$ that a banana, a pack of chips, and a can of Coke are placed on the counter. The LLM planner will first generate three options $\mathcal{C}_{i}$ of bringing each item to the user, and predict the label $\hat{y}_{i}$ corresponding to the can.
Knowledge Base Construction. Consider a training set $\mathcal{U}=\{(x_{1},\mathcal{G}_{1})\}_{i=1}^{N}$ comprising $N$ instances. For each instance, $x_{i}:=(d_{i},o_{i})$ encompasses a pair of the task $d_{i}$ and the observation $o_{i}$, and a set of all valid options $\mathcal{G}_{i}$ satisfying the task specification and the observation. To construct the knowledge base, we query LLM and generate a set of candidate plans $\mathcal{C}_{i}$ with alphabetic label from $\mathcal{Y}$, conditioned on the task $d_{i}$, the observation $o_{i}$, along with hand-crafted few-shot examples. This is followed by prompting the LLM to produce rationale $k_{i}$ given the ground truth valid options $\mathcal{G}_{i}$. Specifically, we use in-context learning with few-shot examples to guide LLM generating explanations of why certain options are valid according to the ground truth. Incorporating ground truth actions directly into the prompt allows LLMs to generate reasoning that more closely aligns with the actual options. To facilitate retrieval during the inference phase, we compute the textual embedding of each instruction $d_{i}$ as the key to the knowledge and store them in the knowledge base dictionary $\mathcal{K}$. We summarize the procedure of knowledge base construction in Algorithm 1.
@@
$\mathcal{U}=\{(x_{1},\mathcal{G}_{1},),\ldots,(x_{N},\mathcal{G}_{N},))\}$$\mathcal{K}\leftarrow\{\}$Initialize Knowledge Base Dictionary each train example $x_{i}=(d_{i},o_{i})$$\mathcal{C}_{i}\leftarrow\text{GenerateChoices}(x_{i})$$k_{i}\leftarrow\text{GenerateKnowledge}(x_{i},\mathcal{C}_{i},\mathcal{G}_{i})$$e_{i}\leftarrow\text{Embed}(d_{i})$$\mathcal{K}[e_{i}]\leftarrow\{x_{i},\mathcal{C}_{i},k_{i},\mathcal{G}_{i}\}$
@@ 3 Introspective Conformal Prediction
Successful human-centered robot autonomy hinges on accurate comprehension of users’ goals—in situations where a user-specified task admits multiple valid interpretations, it is crucial for the robot to detect task ambiguity and solicit further instructions. Directly querying language models for a prediction (even with few-shot in-context learning strategies) falls short of providing clear confidence indicators. This can result in overconfident decisions that clash with user expectations. Conformal prediction addresses this by quantifying the model’s certainty through confidence level calibration. However, its utility is contingent upon the model’s reasoning capabilities, and it may, in extreme cases, produce a wide array of options–including irrelevant or hazardous ones—to uphold high success rates. In this section, We augment introspective planning with conformal prediction to provide a tighter bound on the statistical guarantee of success. The synergy of these approaches is illustrated in Figure 2.
@@
Extending to Multi-Label Prediction. In many cases, task specifications can be inherently ambiguous. This is particularly apparent when there are multiple distinct candidate plans that align with the task, even if the user may only prefer one of them. Consequently, the robot must reason about the degree of certainty regarding the ambiguous nature of the specified task. The objective here is to mitigate under-coverage which could potentially result in deviation from the true intent, as well as to preclude invalid options that could pose safety risks.
In this context, our proposed method, introspective conformal prediction, can be seamlessly adapted to evaluate the planner’s level of uncertainty when predicting the exact set under ambiguous specifications. This process involves extracting prospective options from the power set of candidate plans $P(\hat{\mathcal{C}}_{\text{test}})$ instead of $\mathcal{C}_{\text{test}}$ and then query the language model to obtain joint confidence score associated with this set of options.
@@ 4.1 Evaluation Metrics
Traditional measures such as success rate and help rate inadequately capture the nuances of LLM behavior, particularly in scenarios requiring precision and accuracy. We introduce additional metrics to quantify prediction’s overconfidence in ambiguous tasks and imprecision when the LLM mixes valid and invalid options in their planning.
@@
  • Exact Set Rate (ESR) quantifies the occurrence of predictions precisely matching with all valid actions derived from instructions, serving as a crucial measure of accuracy in ambiguous situations. This mirrors the concept of Precision.
  • Non-compliant Contamination Rate (NCR) calculates the proportion of prediction sets containing options that are invalid for the task, measuring the LLM’s ability to comply with the instruction.
@@ 5 Results
We implemented all tasks using OpenAI’s GPT-3.5 (text-davinci-003) and GPT-4 Turbo (gpt-4-1106-preview). Results on GPT-3.5 are presented in the main paper, with GPT-4 Turbo findings detailed in Appendix C. We employed Sentence-BERT (Reimers & Gurevych, 2019) to encode instructions, retrieving the top $m=3$ text embeddings based on cosine similarity. Our knowledge base mirrors the calibration set in size at 400. However, experimental explorations with sizes of 10, 50, 100, and 200 indicated that a smaller knowledge base, specifically of 100 or 200, effectively achieves satisfactory results. The relationship between performance and knowledge base size is further detailed in Appendix B.
@@ 5.1 Direct Prediction
Results from Table 1 provide unique insights into various predictive models’ performance. In general, we found the proposed introspective reasoning with direct prediction surpasses all baseline methods in the category of Success Rate (SR) and Exact Set Rate (ESR). Retrieve-Q-CoT and Auto-CoT don’t outperform other prompt methods on the original dataset, since the model may potentially generate misleading reasoning in the training phase without access to ground truth actions. However, they excel in the unambiguous instruction dataset, achieving 70% ESR, despite lower success and minimal help rates across all datasets. This observed trend suggests both methods lean towards generating incorrect predictions with high confidence, thereby highlighting the critical role of introspective planning in offering meaningful guidance for enhancing LLM reasoning capabilities. Moreover, as shown in Table 2, our approach significantly outperforms all baseline methods in scenarios that are safety-critical, exhibiting impressive improvements in safety prioritization and dramatically reducing Noncompliance Contamination Rate (NCR) and Unsafe Contamination Rate (UCR).
Through additional conformal prediction, KnowNo is able to achieve a notably higher success rate in comparison to Prompt Set at the sacrifice of low Exact Set Rate (ESR)—indicating the inclusion of incorrect options. In contrast, the direct prediction method is distinguished by its ability to generate more precise and accurate prediction sets, with the Prompt Set method substantially improving the ESR by nearly 19.5% over KnowNo, an enhancement that is further augmented through the strategic integration of Chain of Thought (CoT) techniques. In the following section, we will provide a comprehensive analysis, meticulously highlighting the numerous substantial advantages and benefits that arise from integrating introspective planning approaches with conformal prediction techniques, especially when compared to the traditional prompting methods.
@@ 5.2 Conformal Prediction
We compare introspective reasoning to two baseline methods within the conformal prediction framework, as detailed in section 3, by conducting detailed ablation studies to investigate the impact of Target Success Rate (target success rate) on the model’s performance. The results, illustrated in 3(a) and 7(b), demonstrate that conformal prediction effectively calibrates all three methods, enabling the empirical success rate to align with target success rate across both examined datasets. Furthermore, 4(a) reveals that our proposed introspective reasoning method consistently achieves a higher Exact Set Rate across the full range of target success rate, outperforming the Retrieval-Q-CoT, which only shows marginal improvement over KnowNo at target success rate levels below $85\%$. This trend is similarly observed in the safe mobile manipulation dataset, as depicted in 5(a), reinforcing the effectiveness of our approach.
The observation suggests that Retrieve-Q-CoT and KnowNo exhibit less certainty regarding the correct options, necessitating the inclusion of additional choices in their prediction sets to achieve the desired level of success rate. To validate this observation, we calculated the average prediction set size of each method and have detailed our findings in 4(c) and 5(c). Our analysis revealed that introspective planning tends to include fewer options in its prediction sets for a given target success rate. This is also reflected in each model’s help rate across both datasets, as illustrated in 4(b) and 5(b), where it is evident that both KnowNo and Retrieval-Q-CoT require human clarifications more frequently than introspective planning. From the conformal prediction perspective, we noticed that introspective planning has a higher calibration threshold $1-\hat{q}$ than two baseline methods in both datasets (4(d) and 5(d)), which indicates our methods achieve a tighter confidence bound for the statistical guarantee. These comprehensive empirical results strongly suggest that introspective planning offers a more precise reasoning framework when navigating uncertainties in robot planning scenarios.
Furthermore, we also study whether introspective planning helps to better reason about task ambiguities and unsafe actions in robot planning. Our preliminary findings, drawn from experiments conducted on Safe Mobile Manipulation dataset, demonstrate that introspective planning not only significantly betters the non-compliant contamination rate when compared to the two baseline approaches but also consistently maintains a lower rate of unsafe contamination throughout all target success rate. Finally, we exhibit the capability of extending the current conformal prediction framework to accommodate multi-label prediction settings and assess its performance in Appendix D.
@@ Appendix C Additional Quantitative Results on GPT-4 Turbo
Appendix D Additional Quantitative Results on Multi-Label Conformal Prediction
We conduct multi-label prediction tasks with introspective planning and conformal prediction using GPT-4 Turbo (gpt-4-1106-preview) on the KnowNo Dataset. The task success in multi-label setting requires the model to predict a single multi-label prediction set that matches the ground truth exactly. We introduce two modified metrics in this experiment.
@@ Appendix D Additional Quantitative Results on Multi-Label Conformal Prediction
As depicted in 12(a), our proposed method surpasses two baselines in terms of the overask rate at high target success rates ranging from 0.8-0.95. This suggests that introspective reasoning aids in uncertainty reduction through guided reasoning, thereby minimizing conservativeness. At a lower target success rate, Retrieval-Q-CoT and introspective reasoning demonstrate comparable performance. As the target success rate nears 1, all methods exhibit expected conservatism, necessitated by the requirement to include all options in the prediction.
From 12(b), it can be observed that the Retrieval-Q-CoT method consistently attains a zero overstep rate across the entire target success rate spectrum. Our proposed technique also exhibits low rates compared to KnowNo. It should be noted that as the target success rate escalates, the quantity of confident predictions generated from all approaches diminishes, eventually dwindling to zero.
@@ Appendix F Additional Qualitative Results
You are simulating a robot operating in an office kitchen. You are in front of a counter with two closed drawers, a top one and a bottom one. There is also a landfill bin, a recycling bin, and a compost bin. Please follow the template exactly to generate your response. Scene: On the counter, there is a bottled water, a bottled unsweetened tea, and a Sprite. Task: I want a drink that’s not too carbonated. Options: A) pick up the sprite B) pick up the bottled unsweetened tea C) pick up the bottled water D) do nothing E) an option not listed here Explain: The task specifies a preference for a drink that is ”not too carbonated,” which rules out the Sprite, a carbonated beverage. The bottled water and the bottled unsweetened tea are both non-carbonated options. Therefore, picking up the bottled unsweetened tea and picking up the bottled water align with the given task. Prediction: B, C Scene: On the counter, there is a orange soda, a RedBull, and a Pepsi. Task: Bring me something with caffeine. Options: A) pick up the orange soda B) pick up the orange soda and the pepsi C) pick up the pepsi D) an option not listed here E) pick up the redbull Explain: The task requires selecting a beverage that contains caffeine. Typically, orange soda does not contain caffeine, while Pepsi and RedBull do. Therefore, options C (picking up the Pepsi) and E (picking up the RedBull) are the correct actions that align with the task’s requirement for a caffeinated drink. Option A is incorrect because orange soda usually doesn’t have caffeine, and option B is partially incorrect because it includes the orange soda. If there is another caffeinated beverage not listed, it would fall under option D, but based on the items provided, C and E are the ground truth actions. Prediction: C, E Scene: On the counter, there is a orange soda, a bottled unsweetened tea, and a bottled water. Task: I want a drink, but nothing too sweet. Options: A) an option not listed here B) pick up the bottled unsweetened tea C) pick up the orange soda D) pick up the pepsi E) pick up the bottled water Explain: The task specifies a preference for a drink that is ”not too sweet.” Among the options provided, the orange soda is typically sweet, so it would not meet the criteria. The bottled unsweetened tea and the bottled water are both likely to be less sweet than the soda. Therefore, options B (bottled unsweetened tea) and E (bottled water) align with the request for a less sweet drink. Option D (pick up the Pepsi) is not applicable as Pepsi is not mentioned as being on the counter in the scene. Prediction: B, E Scene: On the counter, there is a sparkling water, a grape soda, and a Mountain Dew. Task: I want a drink with a kick of caffeine and citrus flavor Options: A) pick up the mountain dew B) pick up the sparkling water C) pick up the grape soda D) an option not listed here E) do nothing Explain: The task specifies a preference for a drink with both caffeine and citrus flavor. Mountain Dew is known for its citrus flavor and caffeine content, making it the appropriate choice for this task. Sparkling water is typically not caffeinated and does not have a strong citrus flavor unless it is specifically a citrus-flavored variety, which is not indicated here. Grape soda generally does not contain caffeine and has a grape flavor rather than citrus. Therefore, option A (picking up the Mountain Dew) is the correct action that aligns with the task’s requirements. Options B and C do not meet the criteria of having both caffeine and citrus flavor. Prediction: A

### 18 Feb 2024 v2

@@ 2 Introspective Planning
The fundamental aim of introspective planning is to guide LLMs automatically to reason about task uncertainty and assess the feasibility of candidate actions. While this guidance can take various forms, our implementation here is based on distilling example reasoning processes from a knowledge base to inform the LLM via in-context learning.
Problem Formulation. Similar to (Ren et al., 2023), we cast LLM-based planning as multiple-choice question answering (MCQA). Given a task statement $d_{i}$ and the observation $o_{i}$, the LLM planner first generates a set of candidate plans $\mathcal{C}_{i}$, each assigned a unique letter label from $\mathcal{Y}\coloneqq\{\textrm{`}A\textrm{'},\textrm{`}B\textrm{'},\textrm{`}C\textrm{'},\ldots\}$ ’ , … }. The planner then predicts $\hat{y}_{i}\in\mathcal{Y}$, aiming to match the unknown true user intent $z_{i}$. For example, consider the stated task $d_{i}$ring me that soda” with the observation $o_{i}$ that a banana, a pack of chips, and a can of Coke are placed on the counter. The LLM planner will first generate three options $\mathcal{C}_{i}$ of bringing each item to the user, and predict the label $\hat{y}_{i}$ corresponding to the Coke .
Knowledge Base Construction. Consider a training set $\mathcal{U}=\{(x_{i},\mathcal{G}_{i})\}_{i=1}^{N}$ comprising $N$ instances. For each instance, $x_{i}:=(d_{i},o_{i})$ encompasses a pair of the task $d_{i}$ and the observation $o_{i}$, and a set of all valid options $\mathcal{G}_{i}$ satisfying the task specification and the observation. To construct the knowledge base, we query LLM and generate a set of candidate plans $\mathcal{C}_{i}$ with alphabetic label from $\mathcal{Y}$, conditioned on the task $d_{i}$, the observation $o_{i}$, along with hand-crafted few-shot examples. This is followed by prompting the LLM to produce rationale $k_{i}$ given the ground truth valid options $\mathcal{G}_{i}$. Specifically, we use in-context learning with few-shot examples to guide LLM generating explanations of why certain options are valid according to the ground truth. Incorporating ground truth actions directly into the prompt allows LLMs to generate reasoning that more closely aligns with the actual options. To facilitate retrieval during the inference phase, we compute the textual embedding of each instruction $d_{i}$ as the key to the knowledge and store them in the knowledge base dictionary $\mathcal{K}$. We summarize the procedure of knowledge base construction in Algorithm 1.
@@
$\mathcal{U}=\{(x_{1},\mathcal{G}_{1},),\ldots,(x_{N},\mathcal{G}_{N},)\}$$\mathcal{K}\leftarrow\{\}$Initialize Knowledge Base Dictionary each train example $x_{i}=(d_{i},o_{i})$$\mathcal{C}_{i}\leftarrow\text{GenerateChoices}(x_{i})$$k_{i}\leftarrow\text{GenerateKnowledge}(x_{i},\mathcal{C}_{i},\mathcal{G}_{i})$$e_{i}\leftarrow\text{Embed}(d_{i})$$\mathcal{K}[e_{i}]\leftarrow\{x_{i},\mathcal{C}_{i},k_{i},\mathcal{G}_{i}\}$
@@ 3 Introspective Conformal Prediction
Successful human-centered robot autonomy hinges on accurate comprehension of users’ goals—in situations where a user-specified task admits multiple valid interpretations, it is crucial for the robot to detect task ambiguity and solicit further instructions. Directly querying language models for a prediction (even with few-shot in-context learning strategies) falls short of providing clear confidence indicators. This can result in overconfident decisions that clash with user expectations. On the other hand, conformal prediction offers the advantage of providing quantifiable confidence levels for its predictions, enabling a clearer understanding of a model’s certainty in its outcomes. However, its effectiveness can be compromised if the underlying model lacks strong reasoning abilities. In extreme cases, to maintain high success rates, it might output an excessively broad range of options, including irrelevant or unsafe ones. In this section, We augment introspective planning with conformal prediction to provide a tighter bound on the statistical guarantee of success. The synergy of these approaches is illustrated in Figure 2.
@@
Extending to Multi-Label Prediction. In many cases, the user’s task statements can truly be ambiguous, and this means that there’ll be multiple distinct candidate plans that comply with what the user asked for in different ways, even though the user may only be satisfied with one of them. Therefore the robot needs to be able to reason about how certain it is that the task was specified ambiguously.
[…]
@@ 4.1 Evaluation Metrics
Traditional measures such as success rate and help rate inadequately capture the nuances of LLM behavior, particularly in scenarios requiring precision and accuracy. We introduce additional metrics to capture prediction’s overconfidence in ambiguous tasks and imprecision when the LLM mixes valid and invalid options in their planning.
@@
  • Exact Set Rate (ESR) measures the frequency of the LLM’s predictions perfectly aligning with all valid actions inferred from instructions. It’s crucial for evaluating the model’s ability to generate precise responses.
  • Non-compliant Contamination Rate (NCR) calculates the proportion of prediction sets containing options that deviate from the given instructions, thereby measuring the LLM’s ability to follow directions accurately.
@@ 5 Results
We implemented all tasks using OpenAI’s GPT-3.5 (text-davinci-003) and GPT-4 Turbo (gpt-4-1106-preview). Results on GPT-3.5 are presented in the main paper, with GPT-4 Turbo findings detailed in Appendix C. We employed Sentence-BERT (Reimers & Gurevych, 2019) to encode instructions, retrieving the top $m=3$ text embeddings based on cosine similarity. The size of our knowledge base is the same as the calibration set which is 400. However, experimental explorations with sizes of 10, 50, 100, and 200 indicated that a smaller knowledge base, specifically of 100 or 200, effectively achieves satisfactory results. The relationship between performance and knowledge base size is further detailed in Appendix B.
@@ 5.1 Direct Prediction
In Table 1, our analysis highlights that introspective planning with direct prediction outperforms all baseline methods in Success Rate (SR) and Exact Set Rate (ESR), indicating superior predictive accuracy. Retrieve-Q-CoT and Auto-CoT don’t outperform other prompt methods on the original dataset, since the model may potentially generate misleading reasoning in the training phase without access to ground truth actions. However, they excel in the unambiguous instruction dataset, achieving 70% ESR, despite lower success and minimal help rates across all datasets. This observed trend suggests both methods lean towards generating incorrect predictions with high confidence, thereby highlighting the critical role of introspective planning in offering meaningful guidance for enhancing LLM reasoning capabilities. Moreover, as shown in Table 2, our approach significantly outperforms all baselines in safety-critical scenarios, exhibiting impressive improvements in safety prioritization and dramatically reducing Noncompliance Contamination Rate (NCR) and Unsafe Contamination Rate (UCR).
Through additional conformal prediction, KnowNo is able to achieve a notably higher success rate in comparison to Prompt Set at the sacrifice of low Exact Set Rate (ESR)—indicating the inclusion of incorrect options. In contrast, the direct prediction method is distinguished by its ability to generate more precise and accurate prediction sets, with the Prompt Set method substantially improving the ESR by nearly 19.5% over KnowNo, an enhancement that is further augmented through the strategic integration of Chain of Thought (CoT) techniques. In the next section, we’ll delve into the advantages of integrating introspective planning with conformal prediction, demonstrating their superiority over traditional prompting methods in detail.
@@ 5.2 Conformal Prediction
In section 3, we compare introspective planning with two baselines using conformal prediction, assessing the influence of Target Success Rate (target success rate) through ablation studies. The findings (3(a), 7(b)) confirm that conformal prediction aligns the empirical success rate with target success rate in both datasets. Notably, 4(a) shows that our method consistently achieves a higher Exact Set Rate across the full range of target success rate, outperforming the Retrieval-Q-CoT, which only shows marginal improvement over KnowNo at target success rate levels below $85\%$. This superiority is echoed in the safe manipulation dataset (5(a)), validating our approach’s robustness.
Our analysis indicates Retrieve-Q-CoT and KnowNo are less certain about correct choices, leading to larger prediction sets for achieving desired success rate levels. To validate this observation, we calculated the average prediction set size of each method and have detailed our findings in 4(c) and 5(c). Our analysis revealed that introspective planning tends to include fewer options in its prediction sets for a given target success rate. This is also reflected in each model’s help rate across both datasets, as illustrated in 4(b) and 5(b), where it is evident that both KnowNo and Retrieval-Q-CoT require human clarifications more frequently than introspective planning. From the conformal prediction perspective, we noticed that introspective planning has a higher calibration threshold $1-\hat{q}$ than two baseline methods in both datasets (4(d) and 5(d)), which indicates our methods achieve a tighter confidence bound for the statistical guarantee. These comprehensive empirical results strongly suggest that introspective planning offers a more precise reasoning framework when navigating uncertainties in robot planning scenarios.
Furthermore, we also study whether introspective planning helps to better reason about task ambiguities and unsafe actions in robot planning. Preliminary results from the Safe Mobile Manipulation dataset show introspective planning significantly reduces non-compliant contamination rates compared to baseline methods and maintains a lower unsafe contamination rate across all target success rate levels. Additionally, we demonstrate how the conformal prediction framework can be extended for multi-label predictions, with performance details in Appendix D.
@@ Appendix C Additional Quantitative Results on GPT-4 Turbo
Appendix D Results on Multi-Label Conformal Prediction
We conduct multi-label prediction tasks with introspective planning and conformal prediction using GPT-4 Turbo (gpt-4-1106-preview) on the Mobile Manipulation Dataset. The task is successful if the ground truth set is in the set of prediction sets generated by our multi-label conformal prediction algorithms. We introduce two modified metrics in this experiment. We introduced two modified metrics in this experiment.
@@ Appendix D Results on Multi-Label Conformal Prediction
In the multi-label setting, the robot is certain only if the conformant family of sets only contains one set and that set is a singleton. Conversely, if the family contains multiple sets (even if those sets are singletons) or contains only one set but the set is not a singleton, then the robot is uncertain. In other words, we take the union of all the sets of options in the conformant family, and if that union contains more than one option, we count that as the robot being uncertain.
@@
As depicted in 12(a), our proposed method surpasses two baselines in terms of the overask rate at high target success rates ranging from 0.8-0.95. This suggests that introspective reasoning aids in uncertainty reduction through guided reasoning, thereby minimizing conservativeness. At a lower target success rate, Retrieval-Q-CoT and introspective reasoning demonstrate comparable performance. As the target success rate nears 1, all methods exhibit expected conservatism, necessitated by the requirement to include all options in the prediction. Overall, introspective planning approach performs the best in identifying unambiguous tasks.
From 12(b), it can be observed that the Retrieval-Q-CoT method consistently attains a zero overstep rate across the entire target success rate spectrum. The effectiveness of our method closely mirrors that of Retrieval-Q-CoT and is notably superior to the KnowNo approach. This comparison suggests that our planning method proposes correct choices frequently when it is certain about its prediction.
@@ Appendix F Additional Qualitative Results
Scene: On the counter, there is a bottled water, a bottled unsweetened tea, and a Sprite.
Task: I want a drink that’s not too carbonated.
Options:
A) pick up the sprite
B) pick up the bottled unsweetened tea
C) pick up the bottled water
D) do nothing
E) an option not listed here
Explain: The task specifies a preference for a drink that is ”not too carbonated,” which rules out the Sprite, a carbonated beverage. The bottled water and the bottled unsweetened tea are both non-carbonated options. Therefore, picking up the bottled unsweetened tea and picking up the bottled water align with the given task.
Prediction: B, C
Scene: On the counter, there is a orange soda, a RedBull, and a Pepsi.
Task: Bring me something with caffeine.
Options:
A) pick up the orange soda
B) pick up the orange soda and the pepsi
C) pick up the pepsi
D) an option not listed here
E) pick up the redbull
Explain: The task requires selecting a beverage that contains caffeine. Typically, orange soda does not contain caffeine, while Pepsi and RedBull do. Therefore, options C (picking up the Pepsi) and E (picking up the RedBull) are the correct actions that align with the task’s requirement for a caffeinated drink. Option A is incorrect because orange soda usually doesn’t have caffeine, and option B is partially incorrect because it includes the orange soda. If there is another caffeinated beverage not listed, it would fall under option D, but based on the items provided, C and E are the ground truth actions.
Prediction: C, E
Scene: On the counter, there is a orange soda, a bottled unsweetened tea, and a bottled water.
Task: I want a drink, but nothing too sweet.
Options:
A) an option not listed here
B) pick up the bottled unsweetened tea
C) pick up the orange soda
D) pick up the pepsi
E) pick up the bottled water
Explain: The task specifies a preference for a drink that is ”not too sweet.” Among the options provided, the orange soda is typically sweet, so it would not meet the criteria. The bottled unsweetened tea and the bottled water are both likely to be less sweet than the soda. Therefore, options B (bottled unsweetened tea) and E (bottled water) align with the request for a less sweet drink. Option D (pick up the Pepsi) is not applicable as Pepsi is not mentioned as being on the counter in the scene.
Prediction: B, E
Scene: On the counter, there is a sparkling water, a grape soda, and a Mountain Dew.
Task: I want a drink with a kick of caffeine and citrus flavor
Options:
A) pick up the mountain dew
B) pick up the sparkling water
C) pick up the grape soda
D) an option not listed here
E) do nothing
@@
Explain: The task specifies a preference for a drink with both caffeine and citrus flavor. Mountain Dew is known for its citrus flavor and caffeine content, making it the appropriate choice for this task. Sparkling water is typically not caffeinated and does not have a strong citrus flavor unless it is specifically a citrus-flavored variety, which is not indicated here. Grape soda generally does not contain caffeine and has a grape flavor rather than citrus. Therefore, option A (picking up the Mountain Dew) is the correct action that aligns with the task’s requirements. Options B and C do not meet the criteria of having both caffeine and citrus flavor.
Prediction: A